OpenStudy (anonymous):

Imaginary numbers, what is: \[i^i\] I know that i is defined as: \[i = \sqrt{-1}\] So how does this work out?

OpenStudy (ikram002p):

e^(-π/2)

OpenStudy (anonymous):

@ikram002p Please explain :-). How you got that I know that \(e = 2.7182818\) but how do you get that \(e^{\frac{-\pi}{2}} = i^i\), how do we prove it?

OpenStudy (ikram002p):

e^i theta=cos (theta)+i sin(theta) right ?

OpenStudy (anonymous):

I haven't studied that far yet :o. I guess I will have to learn trigonometry to learn how to do this. Thank you though!

OpenStudy (ikram002p):

oh .. dnt need trigonometry, its the complex num formula from tylor expansions , the rest of it so easy :o

OpenStudy (ikram002p):

just keep up with me ok ?

OpenStudy (anonymous):

I only know Algebra II... Taylor's expansion is calculus.

OpenStudy (ikram002p):

i dnt wanna u to prove tylor or somthing , just know this formula that e^i theta=cos (theta)+i sin(theta) ok ? else is so easy u can note in the second term there is "i sin(theta)" so find theta s.t sin theta gives u 1 ,-1 ok ?

OpenStudy (ikram002p):

so if u note that when sin (theta)=1,-1 cos (theta) =0

OpenStudy (ikram002p):

so just take theta = π/2 e^i π/2=cos (π/2)+i sin(π/2) e^i π/2=0+i(1) e^i π/2=i

OpenStudy (ikram002p):

e^i π/2=i raise to the power i ( e^i π/2)^i=i^i e^(i*i π/2)=i^i.............(i*i=-1) e^-π/2=i^i # done :D

OpenStudy (anonymous):

@ikram002p I'm so sorry, to leave you half way :O my internet disconnected, sorry! I understand what you mean, thanks :D

OpenStudy (ikram002p):

np :D